OpenStudy (anonymous):

Evaluate the integral using integration by parts where possible. ((2x+1)/(e^9x))dx

myininaya (myininaya):

\[\int\limits_{}^{}\frac{2x+1}{e^{9x}} dx=\int\limits_{}^{}(\frac{2x}{e^{9x}}+\frac{1}{e^{9x}}) dx\] \[=\int\limits_{}^{}2xe^{-9x} dx+\int\limits_{}^{}e^{-9x} dx\]

myininaya (myininaya):

Does that form look better to you? Or are you still uncertain on how to evaluate this?

OpenStudy (anonymous):

It's a much better form, so is the final answer the actual final answer or does it need more work to it?

myininaya (myininaya):

You to actually integrate You know use integration by parts to integrate the first integral And then the second integral is tons easier

OpenStudy (anonymous):

Just so I know I'm on the right track...e^-9x integrated is still e^-9x right? Or does that require u substitution.

myininaya (myininaya):

use this for the second part! \[\int\limits_{}^{}e^{c \cdot x} dx=\frac{1}{c} e^{c \cdot x} +C \text{ where c is a constant not equal to zero}\]

myininaya (myininaya):

For the first part... You can do a table

OpenStudy (anonymous):

So 1/-9 e^-9x + C? For the second part. The first part I know how to do.

myininaya (myininaya):

Yes that is all for the second part

OpenStudy (anonymous):

Okay cool, thank you very much!

myininaya (myininaya):

So you know how to do the first part which is what I find weird since you didn't know how to do the second part

myininaya (myininaya):

Are you sure you know how to do the first part?

myininaya (myininaya):

You want to tell me what you got for the answer? I will check.

OpenStudy (anonymous):

I can tell you what I get after I finish it if you'd like.

myininaya (myininaya):

That is fine :)

OpenStudy (anonymous):

Because you told me how to do the second part I can do the first part, which is why I didn't need help with the first part.

myininaya (myininaya):

lol Oh okay. Gotcha!

OpenStudy (anonymous):

I'm not really sure if this is right, but i used the formula I u dv = u*v - I v*dv, where I stands for integral. Here's what I got: (2x)((-1/9)e^-9x)+(1/9)-((1/9e^-9x))*((1/9e^-9x))+C It's not simplified at all I realize.

myininaya (myininaya):

Ok it is hard for me to read that but if you are saying what I got here, then you are correct: \[\int\limits_{}^{}fg'dx=fg-\int\limits_{}^{}f'g dx\] \[f=2x, g'=e^{-9x} => f'=2, g=\frac{-1}{9} e^{-9x}\] So plugging into our formula we get: \[\int\limits_{}^{}2x \cdot e^{-9x} dx=2x \cdot \frac{-1}{9}e^{-9x}-\int\limits_{}^{} 2 \cdot \frac{-1}{9}e^{-9x} dx\] \[=\frac{-2}{9}xe^{-9x}+\frac{2}{9} \cdot \frac{-1}{9}e^{-9x} +C \] \[=\frac{-2}{9}xe^{-9x}-\frac{2}{81}e^{-9x}+C\] Is that what you have for the first part?

myininaya (myininaya):

I think you are off by a constant multiply in your second term

myininaya (myininaya):

multiple*

myininaya (myininaya):

Looks like you are missing the constant multiple 2

myininaya (myininaya):

just for the second term if i read right

OpenStudy (anonymous):

What you have is a simplified version of what I had, which is what I didn't know how to do.

myininaya (myininaya):

wait you have e^{-9x}*e^{-9x} you should only have this factor once not twice for the second term

OpenStudy (anonymous):

I do, I just combined the first and second parts so I could get an entire answer.

myininaya (myininaya):

I'm not sure what you are saying, but for the first part we have \[\frac{-2}{9}xe^{-9x}-\frac{2}{81}e^{-9x}+C_2\] And the second part we have \[\frac{-1}{9}e^{-9x}+C_1\] Do you understand how I got or we got both of these parts?

OpenStudy (anonymous):

Yeah, but to answer the original question, part 1 and part 2 have to be combined. You helped me tremendously!

myininaya (myininaya):

That is correct we need to add the parts

myininaya (myininaya):

\[\frac{-2}{9}xe^{-9x}-\frac{2}{81}e^{-9x}+\frac{-1}{9}e^{-9x}+C\]

OpenStudy (anonymous):

I got the answer right! :) Thanks so much for your help!

myininaya (myininaya):

Ok great! Have a nice day.

OpenStudy (anonymous):

Thanks, you too!